OpenStudy (anonymous):

The Student Council attended a convention. For lunch, 12 out of the 29 members chose milk to drink. a. To the nearest tenth of a percent, what percent of the students chose milk? b. To the nearest tenth of a percent, what percent of the students chose another beverage? (4 points)

hero (hero):

12/29 members got sick

hero (hero):

part/whole = %/100 12/29 = x/100

hero (hero):

Solve for x to find the percent

OpenStudy (anonymous):

a. 40% b. 59%

OpenStudy (anonymous):

Sorry, a is 41%

hero (hero):

29x = 1200 x = 1200/29 x = 41.37